3) Corn is produced under perfectly competitive conditions. Corn farmers have U-shaped, long-run average cost curves that reaNEED HELP PLEASE!
Answer #1

A) long run equilibrium price is the minimum of average total cost. in this case the minimum of average cost is $3 per bushel. therefore the price of corn at the long-run equilibrium will be $3 per bushel. Each corn farm is producing 1000 bushels at this price. Total corn demanded = 2600000-200000*3 = 2,000,000. Number of corn farms in the long run = 2,000,000/1000 = 2000 farms

B) output cannot be increased in the short run which means that total number of farms will still be two thousand and each of them will still be producing 1000 bushels. This implies that total supply will be 2,000,000 bushels. The new market price will be (3,200,000-2,000,000)/200,000 = $6 per bushel. The average cost is still $3 which implies that each farm earns a profit of (6-3)*1000 = $3000.